iosangel:

How would I do this ? Write the equation for function g and describe the transformation that takes function f to function g. f(x)=3x+2 g(x)=f(x)−7

iosangel:

I don't understand what its asking me

iosangel:

Is there like a formula that comes with this?

justjm:

No you're not solving or doing anything. It's just saying that g(x) is f(x) - 7, correct? Based on your knowledge of transformations, what does the -7 do?

iosangel:

We translate it?

iosangel:

down?

justjm:

I'll give a guide to help you out. It is best your memorize your transformations, including stretches and compressions. Let \(f(x)\) be a function of x, and \( g(x) \) be a translation of \( f(x) \) \( g(x) = f(x) + a \) You are moving A units up to get g(x) \( g(x) = f(x) - a \) You are moving A units down to get g(x) \( g(x) = f(x+a) \) You are moving A units LEFT to get g(x) \( g(x) = f(x-a) \) You are moving A units RIGHT to get g(x)

justjm:

Correct down by 7 units